OpenStudy (anonymous):

Inverse functions anyone? How do i find the inverse of f(x) = 4e^(x-2) - 11

OpenStudy (anonymous):

To find the inverse, solve for x in terms of f(x). \[f(x)=4e^{x-2}-11\]\[\frac{f(x)-11}{4}=e^{x-2}\]Take the natural log of both sides:\[\ln(\frac{f(x)-11}{4})=\ln(e^{x-2})\]The natural log "takes away" the exponential\[\ln(\frac{f(x)-11}{4})=x-2\]\[\ln(\frac{f(x)-11}{4})+2=x\] Let me know if you're okay with the reasoning to there.

OpenStudy (anonymous):

ok. i suppose getting to that second step is where i need to work on. where e^x-2 is by itself

OpenStudy (anonymous):

from there i understand completely though

OpenStudy (anonymous):

Okay. So let me expand on that part:

OpenStudy (anonymous):

Subtract 11 from both sides:\[f(x)-11=4e^{x-2}\] Divide both sides by 4: \[\frac{f(x)-11}{4}=e^{x-2}\]

OpenStudy (anonymous):

That part?

OpenStudy (anonymous):

yes. i feel once that parts all sorted and we get to the ln's its simple

OpenStudy (anonymous):

Okidoki. Do you know where to go from there?

OpenStudy (anonymous):

i think... just why did 11 get subtracted and not added

OpenStudy (anonymous):

Oh, whoops. That's because I did it wrong :D

OpenStudy (anonymous):

Plus everywhere instead of minus

OpenStudy (anonymous):

ahh ok that makes a lot more sense now hah. its still + 2 though right .. at the end

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

it makes a lot more sense now. thanks a lot

OpenStudy (anonymous):

Wait wait, not done yet.

OpenStudy (anonymous):

oh . theres more expanding hm

OpenStudy (anonymous):

We should be here:\[\ln(\frac{f(x)+11}{4})+2=x\] Do you know what to do from here?

OpenStudy (anonymous):

yes. this is a property i think

OpenStudy (anonymous):

Nah, it's something else.

OpenStudy (anonymous):

so ln(1/4x+11/4x)+2

OpenStudy (anonymous):

Each way is equally valid. The problem is that we're solving for the inverse function, but we don't have f-inverse anywhere in that equation. All you do now is "switch x and y," with f(x) here playing the part of y.

OpenStudy (anonymous):

\[\ln(\frac{x+11}{4})+2=f^{-1}(x)\]